Middlefield Motors stock is expected to be priced at 89.67 dollars in 1 year. It is expected to pay its next dividend, which is expected to be 2.28 dollars, in 1 year. The stock has a beta of 0.47. The market has an expected return of 12.56 percent, the risk-free rate is 4.32 percent, and the inflation rate is 2.32 percent. What is the current price of Middlefield Motors stock?
Answer #1
As per CAPM
expected return = risk-free rate + beta * (expected return on the market - risk-free rate)
Expected return% = 4.32 + 0.47 * (12.56 - 4.32)
Expected return% = 8.19

Current price = (expected price+dividend)/(1+expected return) = (89.67+2.28)/(1+0.0819)=84.99